# Model predictive control for tracking using artificial references: Fundamentals, recent results and practical implementation

### v1

@@ I Introduction
In 2008, the original stabilizing linear MPC for tracking formulation for tracking piece-wise affine references was intrododuced [5, 6] and in 2018 it was extended to non-linear MPC [7]. Based on these seminal papers, the idea of using an artificial reference has been extended to other classical MPC paradigms, such as tracking periodic references [8, 9, 10, 11, 12, 13, 14], economic MPC [15, 16, 17, 9, 8, 18], robust control [19, 20, 21, 22, 23], stochastic MPC [24, 25], zone-control [26, 27, 28], output-tracking [29], path following and obstacle avoidance [30, 31], or distributed/coordinated control [32, 33, 34, 14, 35].
@@ II Fundamentals of MPC for tracking using an artificial reference: the original formulation
    If $y_{r}$ is changed online, then the terminal set ${\mathcal{X}}_{f}$ must be recomputed as a suitable admissible invariant set containing $x_{r}$, which could be computationally expensive. Additionally, this event could lead to the loss of feasibility of the MPC controller.
@@
where $V_{o}(\cdot):\mathbb{R}^{n_{y}}\to\mathbb{R}$ is the offset cost function, ${\mathcal{X}}_{a}\subseteq\mathbb{R}^{2{n_{x}}+{n_{u}}}$ is the terminal invariant set for tracking, and (6f) requires the artificial reference to be a strictly admissible steady state of the system. We note that (6f) considers strictly admissible steady states to avoid a possible loss of controllability that could result from the presence of active constraints on the artificial reference [5, 55]. However, in practice $\sigma$ can be chosen arbitrarily close to $1$, so the artificial reference can be arbitrarily close to the system constraints.
@@
Let $\hat{{\rm\bf{x}}}\doteq(\hat{x}_{0},\dots,\hat{x}_{N})$, $\hat{{\rm\bf{u}}}\doteq(\hat{u}_{0},\dots,\hat{u}_{N-1})$ be any feasible solution of (6) for the current state $x(t)$. Then, problem (6) is feasible for the successor state $Ax(t)+B\hat{u}_{0}$ for any value of $y_{r}$.
@@
where $T\in\mathbb{S}_{\succ}^{{n_{x}}}$ and $S\in\mathbb{S}_{\succ}^{{n_{u}}}$. This formulation avoids the need to compute a terminal invariant set for tracking ${\mathcal{X}}_{a}$ and does not require the use of a terminal cost $V_{f}(\cdot)$. Additionally, it uses a quadratic offset cost function, leading to a QP problem with a simple structure that can be exploited by first-order optimization algorithms (see Section VI).
@@
An intuitive way of seeing the increase in the domain of attraction that is typically obtained when using an artificial reference is to compare (7) with its classical MPC counterpart, i.e., with (4) taking (4e) as $x_{N}=x_{r}$. In (7) all states $x(0)$ that can reach any setpoint $y\in{\mathcal{Y}}_{s}$ in $N$ steps belong to its feasibility region, and therefore to its domain of attraction (as per Theorems 1 and 2). On the other hand, only states $x(0)$ that can reach the setpoit $y_{r}$ in $N$ steps may belong to the domain of attraction of the classical MPC counterpart of (7). Similar reasoning can be used when comparing the domains of attraction of (4) and (6), as illustrated by the following example.
@@ III Extensions of MPC for tracking
We now present extensions of the original linear MPC for tracking formulation [5, 6] to some of the main MPC paradigms: robust MPC, tracking periodic references and economic MPC. We focus on presenting formulations that provide a light-weight introduction into the use of artificial references within these MPC paradigms. Therefore, we focus on linear MPC formulations that resemble the original MPC for tracking presented in presented in Section II. Section IV will present non-linear extensions of MPC for tracking.
@@ III-A Robust MPC
Under the satisfaction of Assumption 2, the robust MPCT for tracking formulation (9) provides the following recursive feasibility and asymptotic stability guarantees.
@@
    is the set of reachable setpoint considering the tightened constraints ${\mathcal{Z}}_{w}$ (cf. Definition 1), then the output $y(t)$ converges asymptotically to the set $y_{r}\oplus(C+DK)\phi_{K}$. Otherwise, output $y(t)$ asymptotically converges to the set $y_{a}^{\circ}\oplus(C+DK)\phi_{K}$, where $y_{a}^{\circ}$ is the optimal reachable setpoit considering the tightened constraints ${\mathcal{Z}}_{w}$, i.e.,
    $y_{a}^{\circ}=\arg\min_{y\in{\mathcal{Y}}_{w}}\|y_{a}-y_{r}\|_{S}^{2}.$
@@ III-B Tracking periodic references
${\mathcal{Y}}_{s}^{\tau}\doteq\left\{y(\cdot)\in\mathbb{R}^{n_{y}}:\begin{array}[]{@{}l@{}}y(t)=Cx(t)+Du(t)\\ x(t+1)=Cx(t)+Du(t)\\ x(t)=x(t{+}\tau),\,u(t)=u(t{+}\tau)\\ (x(t),u(t))\in\sigma{\mathcal{Z}}\end{array},\;\forall t\right\}.$
@@
for $S\in\mathbb{S}_{\succ}^{{n_{y}}}$, is the offset cost function. Constraints (10e)-(10h) force the artificial reference $({\rm\bf{x_{a}}},{\rm\bf{u_{a}}})$ to describe a reachable periodic trajectory (Definition 3) for the given $\sigma\in[0,1)$, which, once again, is taken arbitrarily close to $1$. Finally, (7f) is a terminal equality constraint that forces the predicted terminal state $x_{N}$ to reach the artificial periodic reference.
@@ III-C Harmonic MPC for tracking
One of the main disadvantages of (10) is that its number of constraints and decision variables grows with the period $\tau\in\mathbb{N}$ of the reference, since the system dynamics and constraints must be imposed along the complete period of the artificial reference, as shown in (10e)-(10f). This can result in long solve times if the period is too long. This issue inspired the development of the MPC for tracking formulation presented in [39, 40], named Harmonic MPC (HMPC). This formulation was originally developed for tracking piece-wise affine setpopints, although it has recently been extended to the problem of tracking harmonic (aka sinusoidal) reference trajectories [41]; a particular class of periodic reference that arises in many practical applications. We now present its original setpoint tracking formulation, since it provides a less notation-heavy introduction of the main idea behind it.
@@
for some $\underline{y},\overline{y}\in\mathbb{R}^{n_{y}}$ satisfying, $\underline{y}<0<\overline{y}$. As in the MPC for tracking formulation (7), HMPC considers the problem of tracking a given setpoint $(x_{r},u_{r})$ and makes use of a terminal equality constraint. However, its artificial reference is given by a harmonic signal, whose value at each prediction step $k$ is given by
@@ IV-A Setpoint tracking
${\mathcal{Y}}_{s}$$\doteq\left\{h(x_{a},u_{a}):(x_{a},u_{a})\in{\mathcal{Z}}_{s}\right\}.$ (17b)
@@
Next, we discuss the role and constructive satisfaction or relaxation of the conditions in Assumptions 56.
@@ IV-B Tracking periodic references
The main idea is that the computationally intensive long horizon planner does not need to be re-optimized at every time $t$, while the trajectory tracking MPC requires a fast sampling rate to react to disturbances. The challenge is that running the planner and tracker independently could lead to references that cause feasibility issues in the terminal set constraint of the tracking MPC. This issue is addressed by adding additional constraints to the planning problem that consider the current tracking error on the terminal set, such that the usual candidate solution remains feasible in the tracker, see [12, Sec. 3.4] for details on a tractable formulation. As a result, a standard trajectory tracking MPC can be implemented at each time $t$, while a partially coupled planner can solve a larger optimization problem over a user chosen sampling period. The partial coupling ensures that both optimization problems remain recursively feasible. Furthermore, if a suitable terminal set is designed, then the reference ${\rm\bf{x_{a}}}$ converges to the optimal target in finite time and the state $x$ converges exponentially to the corresponding optimal trajectory ${\rm\bf{x_{a}^{\circ}}}$ [12, Prop. 13].
@@ IV-C Robust tracking MPC
Conditions (23a)–(23b) ensures that the feedback $\kappa$ drives the state $x$ exponentially to a neighborhood around any nominal trajectory $z$, which is called incremental stability. Inequalities (23c) follows naturally if $g_{j},\kappa$ are Lipschitz continuous, see [76, Prop. 5] for analytical formulas. General incremental Lyapunov functions $V_{\delta}$ and feedbacks $\kappa$ can be constructed offline using (control) contraction metrics [76, 77]. [^1] Following [76, Prop. 2,5], condition (23b) does in fact not hold for all $x\in\mathbb{R}^{{n_{x}}}$ but for all states visited during closed-loop operation due to the tightened constraints. In case of only mildly non-linear dynamics, a simple weighted norm $V_{\delta}(x,z)=\|x-z\|_{M}$ and linear feedback $\kappa(x,z,v)=v+K(x-z)$ can be constructed, in which case $\bar{w},\rho$ can be computed as (weighted) Lipschitz constants. Condition (23d) follows naturally if $V_{\delta}$ is defined as a weighted norm or with a contraction metric.
@@ V Application to Learning-based MPC
In many applications, the system model (16) is subject to significant uncertainty and hence online generated data can be used to improve the model knowledge and thus boost performance. However, the fact that our model changes during runtime also poses challenges in determining an optimal steady-state that should be regulated. This issue can be naturally addressed with MPC for tracking formulations as the desired steady-state is jointly optimized in the MPC. Corresponding results for different data-driven and learning-based models can be found in [42, 22, 76, 79, 80].
@@ VI-A Implementing linear MPC for tracking
Computation of the maximal invariant set can be very expensive. Several alternatives are typically considered: use a terminal equality constraint, as in (7); dynamically scale the terminal set [91]; avoid the use of a terminal set by choosing a sufficiently large prediction horizon [7]; or compute an elliptic terminal invariant set [92, §4.1] by solving an optimization problem with linear matrix inequality constraints [93]. The downside of using an ellipsoid is that the resulting optimization problem is no longer a QP, but instead a SOCP/QCQP. However, as previously discussed, there are several state-of-the-art solvers well suited for these problems [86, 87, 94].
@@ VI-B Exploiting the semi-banded structure of linear MPC for tracking
The stage cost $\ell(x_{k},u_{k},x_{a},u_{a})\doteq\|x_{k}-x_{a}\|^{2}_{Q}+\|u_{k}-u_{a}\|^{2}_{R}$, commonly used in many MPC for tracking formulations, includes cross products between $(x_{k},u_{k})$ and $(x_{a},u_{a})$ that break down the strictly banded structure of the optimization problem related to standard MPC [46, 95]. Consider, for example, the cost of the linear MPC tracking formulation with equality constraints (7):
@@
We notice that $H_{B}$ is a banded matrix and $H_{LR}$ is a low-rank matrix. Thus, the matrix defining the quadratic terms of $V(z)$ has a semi-banded structure that prevents the direct use of the efficient optimization techniques developed for the fully banded case of standard MPC [96, 95]. Two different approaches have been presented in the literature to address this issue. The first approach involves using an extended ADMM formulation that recovers the banded structure by considering three primal decision variables in the ADMM formulation [46]. The second approach relies on the use of the Woodbury matrix identity [97], which allows for solving, in an efficient way, the semi-banded systems of equations required to implement the standard ADMM algorithm (instead of the extended one) [48, 45]. The obtained computational times are similar to the ones corresponding to the standard MPC.

### v2

@@ I Introduction
In 2008, the original stabilizing linear MPC for tracking formulation for tracking piece-wise affine references was introduced [5, 6], and in 2018 it was extended to non-linear MPC [7]. Based on these seminal papers, the idea of using an artificial reference has been extended to other classical MPC paradigms, such as tracking periodic references [8, 9, 10, 11, 12, 13, 14], economic MPC [15, 16, 17, 9, 8, 18], robust control [19, 20, 21, 22, 23], stochastic MPC [24, 25], zone-control [26, 27, 28], output-tracking [29], path following and obstacle avoidance [30, 31], or distributed/coordinated control [32, 33, 34, 14, 35].
@@ II Fundamentals of MPC for tracking using an artificial reference: the original formulation
    If $y_{r}$ is changed online, then the terminal set ${\mathcal{X}}_{f}$ must be recomputed as a suitable admissible invariant set containing the new steady-state $x_{r}$, which could be computationally expensive. Additionally, this could lead to the loss of feasibility of the MPC controller.
@@
where $V_{o}(\cdot):\mathbb{R}^{n_{y}}\to\mathbb{R}$ is the offset cost function, ${\mathcal{X}}_{a}\subseteq\mathbb{R}^{2{n_{x}}+{n_{u}}}$ is the terminal invariant set for tracking [5, §2.2], and (6f) requires the artificial reference to be a strictly admissible steady state of the system. The intuition behind this formulation is to include an artificial reference that acts as a proxy of the desired reference. The formulation steers the system state towards the artificial reference while also penalizing the distance between this artificial reference and the desired one. The idea is for both the system state and artificial reference to converge towards the desired reference as $t$ increases.
We note that (6f) considers strictly admissible steady states to avoid a possible loss of controllability that could result from the presence of active constraints on the artificial reference [5, 55]. However, in practice $\sigma$ can be chosen arbitrarily close to $1$, so the artificial reference can be arbitrarily close to the system constraints.
@@
Let $\hat{{\rm\bf{x}}}\doteq(\hat{x}_{0},\dots,\hat{x}_{N})$, $\hat{{\rm\bf{u}}}\doteq(\hat{u}_{0},\dots,\hat{u}_{N-1})$ be any feasible solution of (6) for the current state $x(t)$. Then, problem (6) is feasible for the successor state $Ax(t)+B\hat{u}_{0}$ for any value of $y_{r}\in\mathbb{R}^{{n_{y}}}$.
@@
where $T\in\mathbb{S}_{\succ}^{{n_{x}}}$ and $S\in\mathbb{S}_{\succ}^{{n_{u}}}$ model a quadratic offset cost $V_{o}(\cdot)$. This formulation avoids the need to compute a terminal invariant set for tracking ${\mathcal{X}}_{a}$ and does not require the use of a terminal cost $V_{f}(\cdot)$. Additionally, it uses a quadratic offset cost function, leading to a QP problem with a simple structure that can be exploited by first-order optimization algorithms (see Section VI).
@@
An intuitive way of seeing the increase in the domain of attraction that is typically obtained when using an artificial reference is to compare (7) with its classical MPC counterpart, i.e., with (4) taking (4e) as $x_{N}=x_{r}$. In (7) all states $x(0)$ that can reach any setpoint $y\in{\mathcal{Y}}_{s}$ in $N$ steps belong to its feasibility region, and therefore to its domain of attraction (as per Theorems 1 and 2). On the other hand, only states $x(0)$ that can reach the setpoint $y_{r}$ in $N$ steps may belong to the domain of attraction of the classical MPC counterpart of (7). Similar reasoning can be used when comparing the domains of attraction of (4) and (6), as illustrated by the following example.
@@ III Extensions of MPC for tracking
We now present extensions of the original linear MPC for tracking formulation [5, 6], to some of the main MPC paradigms: robust MPC, tracking periodic references, and economic MPC. We focus on presenting formulations that provide a light-weight introduction into the use of artificial references within these MPC paradigms. Therefore, we focus on linear MPC formulations that resemble the original MPC for tracking presented in Section II. Section IV will present non-linear extensions of MPC for tracking.
@@ III-A Robust MPC
Under the satisfaction of Assumption 2, the robust MPC for tracking formulation (9) provides the following recursive feasibility and asymptotic stability guarantees.
@@
    is the set of reachable setpoint considering the tightened constraints ${\mathcal{Z}}_{w}$ (cf. Definition 1), then the output $y(t)$ converges asymptotically to the set $y_{r}\oplus(C+DK)\phi_{K}$. Otherwise, output $y(t)$ asymptotically converges to the set $y_{a}^{\circ}\oplus(C+DK)\phi_{K}$, where $y_{a}^{\circ}$ is the optimal reachable setpoint considering the tightened constraints ${\mathcal{Z}}_{w}$, i.e.,
    $y_{a}^{\circ}=\arg\min_{y\in{\mathcal{Y}}_{w}}\|y-y_{r}\|_{S}^{2}.$
@@ III-B Tracking periodic references
${\mathcal{Y}}_{s}^{\tau}\doteq\left\{y(\cdot)\in\mathbb{R}^{n_{y}}:\begin{array}[]{@{}l@{}}y(t)=Cx(t)+Du(t)\\ x(t+1)=Ax(t)+Bu(t)\\ x(t)=x(t{+}\tau),\,u(t)=u(t{+}\tau)\\ (x(t),u(t))\in\sigma{\mathcal{Z}}\end{array},\;\forall t\right\}.$
@@
for $S\in\mathbb{S}_{\succ}^{{n_{y}}}$, is the offset cost function. Constraints (10e)-(10h) force the artificial reference $({\rm\bf{x_{a}}},{\rm\bf{u_{a}}})$ to describe a reachable periodic trajectory (Definition 3) for the given $\sigma\in[0,1)$, which, once again, is taken arbitrarily close to $1$. Finally, (10i) is a terminal equality constraint that forces the predicted terminal state $x_{N}$ to reach the artificial periodic reference.
@@ III-C Harmonic MPC for tracking
One of the main disadvantages of (10) is that its number of constraints and decision variables grows with the period $\tau\in\mathbb{N}$ of the reference, since the system dynamics and constraints must be imposed along the complete period of the artificial reference, as shown in (10e)-(10f). This can result in long solve times if the period is too long. This issue inspired the development of the MPC for tracking formulation presented in [39, 40], named Harmonic MPC (HMPC). This formulation was originally developed for tracking piece-wise affine setpoints, although it has recently been extended to the problem of tracking harmonic (aka sinusoidal) reference trajectories [41]; a particular class of periodic reference that arises in many practical applications. We now present its original setpoint tracking formulation, since it provides a less notation-heavy introduction of the main idea behind it.
@@
for some $\underline{y},\overline{y}\in\mathbb{R}^{n_{y}}$ satisfying $\underline{y}<0<\overline{y}$. As in the MPC for tracking formulation (7), HMPC considers the problem of tracking a given setpoint $(x_{r},u_{r})$ and makes use of a terminal equality constraint. However, its artificial reference is given by a harmonic signal, whose value at each prediction step $k$ is given by
@@ IV-A Setpoint tracking
${\mathcal{Y}}_{s}$$\doteq\left\{h(x_{a},u_{a}):(x_{a},u_{a})\in{\mathcal{Z}}_{s}\right\},$ (17b)
@@
Next, we discuss the role and constructive satisfaction or relaxation of the conditions in Assumptions 56. Note that in the absence of Assumption 6, the closed-loop system may in general converge to some steady-state $x_{a}$, which is not the global minimizer of (18).
@@ IV-B Tracking periodic references
The main idea is that the computationally intensive long horizon planner does not need to be re-optimized at every time $t$, while the trajectory tracking MPC requires a fast sampling rate to react to disturbances. The challenge is that running the planner and tracker independently could lead to references that cause feasibility issues in the terminal set constraint of the tracking MPC. This issue is addressed by adding additional constraints to the planning problem that consider the current tracking error on the terminal set, such that the usual candidate solution remains feasible in the tracker, see [12, Sec. 3.4] for details on a tractable formulation. As a result, a standard trajectory tracking MPC can be implemented at each time $t$, while a partially coupled planner can solve a larger optimization problem over a user chosen sampling period. The partial coupling ensures that both optimization problems remain recursively feasible. Furthermore, if a suitable terminal set is designed, then the reference ${\rm\bf{x_{a}}}$ converges to the optimal target in finite time and the state $x$ converges exponentially to the corresponding optimal trajectory ${\rm\bf{x_{a}^{\circ}}}$ [12, Prop. 13]. This strategy has been further refined in [75] to address (non-periodic) motion-planning problems in embedded autonomous mobile robots.
@@ IV-C Robust tracking MPC
Conditions (23a)–(23b) ensures that the feedback $\kappa$ drives the state $x$ exponentially to a neighborhood around any nominal trajectory $z$, which is called incremental stability. Inequalities (23c) follows naturally if $g_{j},\kappa$ are Lipschitz continuous, see [77, Prop. 5] for analytical formulas. General incremental Lyapunov functions $V_{\delta}$ and feedbacks $\kappa$ can be constructed offline using (control) contraction metrics [77, 78]. [^1] Following [77, Prop. 2,5], condition (23b) does in fact not hold for all $x\in\mathbb{R}^{{n_{x}}}$ but for all states visited during closed-loop operation due to the tightened constraints. In case of only mildly non-linear dynamics, a simple weighted norm $V_{\delta}(x,z)=\|x-z\|_{M}$ and linear feedback $\kappa(x,z,v)=v+K(x-z)$ can be constructed, in which case $\bar{w},\rho$ can be computed as (weighted) Lipschitz constants. Condition (23d) follows naturally if $V_{\delta}$ is defined as a weighted norm or with a contraction metric.
@@ V Application to Learning-based MPC
In many applications, the system model (16) is subject to significant uncertainty and hence online generated data can be used to improve the model knowledge and thus boost performance. However, the fact that our model changes during runtime also poses challenges in determining an optimal steady-state that should be regulated. This issue can be naturally addressed with MPC for tracking formulations as the desired steady-state is jointly optimized in the MPC. Corresponding results for different data-driven and learning-based models can be found in [42, 80, 22, 77, 81, 82].
@@ VI-A Implementing linear MPC for tracking
Computation of the maximal invariant set can be very expensive. Several alternatives are typically considered: use a terminal equality constraint, as in (7); dynamically scale the terminal set [93]; avoid the use of a terminal set by choosing a sufficiently large prediction horizon [7]; or compute an elliptic terminal invariant set [94, §4.1] by solving an optimization problem with linear matrix inequality constraints [95]. The downside of using an ellipsoid is that the resulting optimization problem is no longer a QP, but instead a SOCP/QCQP. However, as previously discussed, there are several state-of-the-art solvers well suited for these problems [88, 89, 96].
@@ VI-B Exploiting the semi-banded structure of linear MPC for tracking
The stage cost $\ell(x_{k},u_{k},x_{a},u_{a})\doteq\|x_{k}-x_{a}\|^{2}_{Q}+\|u_{k}-u_{a}\|^{2}_{R}$, commonly used in many MPC for tracking formulations, includes cross products between $(x_{k},u_{k})$ and $(x_{a},u_{a})$ that break down the strictly banded structure of the optimization problem related to standard MPC [46, 97]. Consider, for example, the cost of the linear MPC for tracking formulation with equality constraints (7):
@@
We notice that $H_{B}$ is a banded matrix and $H_{LR}$ is a low-rank matrix. Thus, the matrix defining the quadratic terms of $V(z)$ has a semi-banded structure that prevents the direct use of the efficient optimization techniques developed for the fully banded case of standard MPC [98, 97]. Two different approaches have been presented in the literature to address this issue. The first approach involves using an extended ADMM formulation that recovers the banded structure by considering three primal decision variables in the ADMM formulation [46]. The second approach relies on the use of the Woodbury matrix identity [99], which allows for solving, in an efficient way, the semi-banded systems of equations required to implement the standard ADMM algorithm (instead of the extended one) [48, 45]. The obtained computational times are similar to the ones corresponding to standard MPC.